OpenStudy (anonymous):

every month, john buys $10.00 of paper per month for his printer. Originally the price was $0.005 per sheet, but the price per sheet increases at a constant rate of $0.0001 per year. How long will it take John to buy his first 500,000 sheets of paper? Set up the differential equation for above problem

OpenStudy (dumbcow):

\[P(t) = .005+\frac{.0001}{12}t\] where t is num of months, p is price of sheet \[N(t) = \frac{10}{P(t)}\] N is num of sheets purchased with $10 at month t \[500,000 = \sum_{t=0}^{T} N(t)\] find T where does the diff equ come in?

OpenStudy (anonymous):

@dumbcow i agree with you. i was thinking of the same.

OpenStudy (anonymous):

because i am taking differential equation class, and this is one of the first order differential equation modeling problem

OpenStudy (anonymous):

I guess that we consider the $10.00 per month is the initial condition

OpenStudy (dumbcow):

is he buying only once a month or do you assume hes buying continuously usually a diif equ model deals with continuous change

OpenStudy (anonymous):

continuous as \[t \rightarrow \]

OpenStudy (anonymous):

\[t \rightarrow \infty \]

OpenStudy (anonymous):

differential equation can be used to understand that there is a linear increase of the cost of the sheets per year.

OpenStudy (anonymous):

make sense

OpenStudy (anonymous):

sorry?

OpenStudy (anonymous):

i mean if we can write the DE as function of time from the amount of sheets of paper with given IVP (initial value problem) to find C constant.

OpenStudy (anonymous):

then with the condition of 500,000 sheets of paper, we can use that equation to find out how long it take for john to buy it.

OpenStudy (anonymous):

look, cost of the sheet.=0.0001+[y/12], where y is the no. of months. [.] denotes the integral part.

OpenStudy (anonymous):

\[500,000=10.12(\frac{ 1 }{ 0.005 }+\frac{ 1 }{ 0.0051 }+..)\] now, if you can find the sum of the harmonic progession, then you can find the time taken.

OpenStudy (anonymous):

awesome, thank you guys

OpenStudy (anonymous):

you're welcome! :)